# Measurement-induced phase transition in periodically driven free-fermionic systems

### v2

@@ V Conclusions
In this work, we investigate the effect of a periodic drive in a continuously monitored free-fermionic lattice model. We find that in the high-frequency limit $T<1$, the results are almost indistinguishable to the no-drive results for the effective hopping strength $J_{eff}=\epsilon/2$ at least for our accessible system size. With the growing belief of the absence of measurement-induced phase transition in the no-drive scenario Starchl et al. (2024b); Müller et al. (2025), it is tempting to conclude that in the high-frequency limit, maybe we don’t have any measurement-induced transition in the thermodynamic limit. No-drive results are presented in the Appendix D for our protocol. It is essential to note that the numerical data are not conclusive enough to confirm the absence of measurement-induced transitions in non-interacting lattice systems like ours. Moreover, the continuum models (which were primarily used as an analytical tool to predict the absence of transition) of Dirac fermions do not guarantee to describe the dynamics of monitored fermions on a tight-binding lattice, mostly because of the lack of energy conservation in such non-unitary dynamics Müller et al. (2025). On the other hand, we find that increasing $T$ tends to promote entanglement growth, and for reasonably large $T$ (in our manuscript, we primarily focus on $T=5$), the data suggest that the measurement-induced transition belonging to the BKT universality class may prevail even in the thermodynamic limit. However, in the absence of any concrete theoretical prediction, it is almost impossible to rule out the possibility that for extremely large system sizes, the steady-state entanglement data will keep showing area law scaling, and whatever we are observing in numerics is just a finite-size crossover between the logarithmic-area law phase. Even in this case, it is safe to say that with increasing $T$, the finite size crossover length scale will certainly increase, and for any reasonably large $T$, such system sizes will be almost inaccessible for any practical purposes Koh et al. (2023a). We also investigate a situation where we take a reasonably large $\gamma$ for which the $T=5$ data of the effective local central charge shows bending in the large subsystem size regime, and then we increase $T$. As expected, we could reduce the bending rate with increasing $T$, but were unable to achieve a scenario where the effective local central charge remains constant. Hence, the analysis does not allow us to conclude whether there is a finite critical driving period $T^{*}$ above which the bending completely stops. The absence of bending in the effective central charge data implies a critical phase. Hence, we also can’t rule out the possibility that beyond a specific critical measurement strength, the system will always remain in the area law phase for any finite driving period.

### v3

@@ V Conclusions
In this work, we investigate the effect of a periodic drive in a continuously monitored free-fermionic lattice model. We find that in the high-frequency limit $T<1$, the results are almost indistinguishable to the no-drive results for the effective hopping strength $J_{eff}=\epsilon/2$ at least for our accessible system size. With the growing belief of the absence of measurement-induced phase transition in the no-drive scenario Starchl et al. (2024b); Müller et al. (2025), it is tempting to conclude that in the high-frequency limit, maybe we don’t have any measurement-induced transition in the thermodynamic limit. No-drive results are presented in the Appendix D for our protocol. It is essential to note that the numerical data are not conclusive enough to confirm the absence of measurement-induced transitions in non-interacting lattice systems like ours, probably because of the exponentially diverging correlation length predicted from NLSM. On the other hand, we find that increasing $T$ tends to promote entanglement growth, and for reasonably large $T$ (in our manuscript, we primarily focus on $T=5$), the data suggest that the measurement-induced transition belonging to the BKT universality class may prevail even in the thermodynamic limit. However, in the absence of any concrete theoretical prediction, it is almost impossible to rule out the possibility that for extremely large system sizes, the steady-state entanglement data will keep showing area law scaling, and whatever we are observing in numerics is just a finite-size crossover between the logarithmic-area law phase. Even in this case, it is safe to say that with increasing $T$, the finite size crossover length scale will certainly increase, and for any reasonably large $T$, such system sizes will be almost inaccessible for any practical purposes Koh et al. (2023a). We also investigate a situation where we take a reasonably large $\gamma$ for which the $T=5$ data of the effective local central charge shows bending in the large subsystem size regime, and then we increase $T$. As expected, we could reduce the bending rate with increasing $T$, but were unable to achieve a scenario where the effective local central charge remains constant. Hence, the analysis does not allow us to conclude whether there is a finite critical driving period $T^{*}$ above which the bending completely stops. The absence of bending in the effective central charge data implies a critical phase. Hence, we also can’t rule out the possibility that beyond a specific critical measurement strength, the system will always remain in the area law phase for any finite driving period.
@@ Appendix F Fluctuation of steady state entanglement entropy
Appendix G Correlation function analysis for $T=5$
In addition to the entanglement entropy, in the case of $T=5$, here we further calculate the equal time connected correlation function in steady state, which is defined as,
$C^{corr}(l)\equiv\overline{|D_{{l+i},i}|^{2}}=\overline{\langle\hat{n}_{l+i}\rangle\langle\hat{n}_{i}\rangle}-\overline{\langle\hat{n}_{l+i}\hat{n}_{i}\rangle}.$ (22)
Here, the overline denotes averaging over different trajectories as well as over positions ($i$) in the steady state. In the critical phase, the correlation function is expected to decay algebraically, whereas in the area-law phase, an exponential decay of the connected correlation function is anticipated Alberton et al. (2021). For both projective and jump measurement protocols, deviations from a power-law dependence in the small $\gamma$ regime have been numerically reported in the absence of drive Poboiko et al. (2023); Starchl et al. (2024a).
In our analysis, we compute the connected steady-state correlation function (see Fig. 24) for $T=5$. We observe that for $\gamma<\gamma_{c}=0.05\pm 0.01$ (assuming a BKT transition exists), particularly at $\gamma=0.03$ and $0.02$, there are no significant signs of deviation from algebraic decay within the system sizes accessible to us. This observation aligns with the entanglement entropy data. However, we cannot rule out the possibility that such deviations might emerge at much larger length scales.
Figure 24: Equal time connected correlation function at steady state $\overline{C^{corr}}$ vs. $\tilde{l}$ plot for $T=5$. The black dashed line represents the fitting function $\overline{C^{corr}}\sim\tilde{l}^{-2}$, where $\tilde{l}=\frac{L}{\pi}\sin(\pi l/L)$. The maximum system size is taken as $L=1004$. Inset shows the variation of $|d\ln\overline{C^{corr}}/d\ln\tilde{l}|$ vs $\tilde{l}$, which seems to saturate to $2$ for $\gamma=0.03$, $0.02$ for large $\tilde{l}$.